# RangeLDM: Fast Realistic LiDAR Point Cloud Generation

### v1

@@ 4.4 Implementation Details
Circular Convolution.
Range images are inherently circular, which means that the left boundary of a range image is actually connected to its right boundary. However, standard convolutions employ zero-padding or symmetry padding and do not take into account such constraints. Thus, as in LiDARGen [94], we replace standard convolutions in both the VAE and LDM with circular convolutions [58], which treat the left and right boundaries as connected neighbors in the topology. Since the 2D convolution operator satisfies translation invariance, it is evident that the entire network is invariant to horizontal shifting of the range image (i.e., rotation in the xy-plane of the point cloud).
@@
Conditioning the Model with Direction.
Real LiDAR point clouds collected from driving scenes usually have a certain directionality. For example, since vehicles typically follow the direction of the road, the x-axis direction of the LiDAR point cloud usually aligns with the road direction. However, since the whole network of our model is invariant to rotation in the xy-plane of the point cloud, the point clouds generated from random noise could result in arbitrary directions. To avoid this, we generate point clouds with directional conditioning. Specifically, we use an $h\times w$ matrix as the condition, in which only the values from the first column are set to $1$, while the rest are all set to 00. This $\mathbf{1}$-valued column corresponds to the x-axis of the point cloud. Such simple conditioning demonstrates effective control over the direction of the generated point cloud, as presented in the qualitative results of the supplementary material.
Network Architectures and Model Hyperparameters.
For simplicity, we adopted a similar architecture to that in [54]. The downsampling factor $f$ is set as $4$ to strike a good balance between efficiency and realistic results. The VAE comprises three encoder blocks and three decoder blocks, with channel settings of 64, 128, 256, 256, 128, and 64. The LDM consists of four downsampling blocks with the last three followed by a transformer layer, and four upsampling blocks with the first three blocks also followed by a transformer layer. The channel settings for the blocks are as follows: 128, 128, 256, 256, 256, 256, 128, 128.
During training, the VAE was optimized using the Adam optimizer with a learning rate of 4.5e-6, while the LDM was trained with the AdamW optimizer with a learning rate of 1e-4. We configured the diffusion steps to be 1000. In the generation process, we employed 50 denoising steps with Denoising Diffusion Implicit Models (DDIM) sampler [65] for point cloud generation. The code of this work will be made public upon the acceptance of the paper, where more implementation details can be found.
@@ 5.1 Experimental Setup
Datasets.
We evaluate our model on two challenging datasets, KITTI-360 and nuScenes. KITTI-360 has 81,106 LiDAR readings from 9 sequences in Germany, covering diverse scenes. We used the first two sequences for validation and trained on the rest. NuScenes is a public dataset with 297,737 LiDAR sweeps in the training set and 52,423 in the testing set, collected in Boston and Singapore.
Method Years $\text{MMD}_{\text{BEV}}$downarrow $\text{FRD}_{\text{range}}$downarrow $\text{JSD}_{\text{BEV}}$downarrow
@@
Evaluation metrics.
Following [94], , we employed three metrics to perform quantitative analysis: Maximum Mean Discrepancy (MMD), Jensen-Shannon divergence (JSD), and Frechet Range Distance (FRD score). We use a $100\times 100$ 2D histogram on the BEV plane to calculate MMD and JSD metrics. The Frechet Range Distance (FRD) score [94] is a metric used to evaluate the quality of samples acquired by a generative model, inspired by the FID score for images [24]. To compute the FRD score, we use RangeNet++ [45], an encoder-decoder-based network for segmentation, which is pre-trained on KITTI-360.
Baselines.
We evaluated our approach for LiDAR point cloud generation against several competitive methods. Specifically, we compared our results with LiDAR VAE [5] and UltraLiDAR [78], both VAE-based methods, LiDAR GAN [5] and Projected GAN [57], which utilize GANs, and LiDARGen [94], an energy-based model.
5.2 Unconditional Generation
5.2.1 Evaluation on the KITTI-360 Dataset
Quantitative Results.
Table 2 displays the quantitative results of the proposed RangeLDM and competing algorithms. The results clearly illustrate that our method significantly outperforms the baselines across all metrics. Notably, our results on the MMD metric are an order of magnitude lower than those of other methods. This underscores the remarkable expressive capabilities of our model in the context of LiDAR point cloud generation.
Table 3: Human study results on KITTI-360.
@@ 5.2.1 Evaluation on the KITTI-360 Dataset
Method Throughput $\uparrow$ (samples/s)
LiDARGen [94] 0.02
$\text{UltraLiDAR}^{\dagger}$ [78] 0.16
Ours 4.86
Table 3: Human study results on KITTI-360.
Table 4: Inference speed on a single RTX 3090 GPU. We generated 1000 samples for calculating the throughput. †: Reproduced via https://github.com/myc634/UltraLiDAR_nusc_waymo.
Qualitative Results.
Figure 4 displays a set of randomly generated samples from competing algorithms, alongside real point cloud samples extracted from the KITTI-360 dataset for comparison. We observe that the Projected GAN [57] exhibits noticeable artifacts in the distant range. LiDARGen [94] effectively captures the general layout, but produces rather noisy outputs and lacks the same degree of straight walls as actual samples. UltraLiDAR [78] generates structured and reasonable scenes, yet it only creates voxelized point clouds (as magnified in the first row) without intensity features. In contrast, our model consistently surpasses the baseline models and yields results closely resembling real-world data. For example, we can generate cars (emphasized by red boxes in the first row of Figure 4) and road bollards (red boxes in the second row of Figure 4) on the road. Additionally, our model generates various scenes such as straightway (the first two rows of Figure 4) and crossroads (the last row of Figure 4). For further unconditional generation results over KITTI-360, please refer to the supplementary materials.
Human study.
To assess the perceptual quality, we conducted an A/B test involving a group of 14 researchers with LiDAR expertise. Following the same evaluation system as [94] and [78], we present pairs of randomly selected images from two point clouds and ask participants to determine which one appeared more realistic. The results, as displayed in Table 4, unequivocally demonstrate the superior visual quality of generation results by our model. In most of the cases, testers favored our results over the baselines.
Model Size.
The number of parameters in LiDARGen [94] and UltraLiDAR [78] are 29.7M and 40.3M, respectively. To ensure fair competition, we limited the capacity of our model to be similar to other approaches. Our model comprises 41.4M parameters, consisting of 12.7M parameters for VAE and 28.7M parameters for LDM. As illustrated in Table 2, our model outperforms the baselines by a large margin under similar model sizes.
Generation Efficiency.
The proposed dimensionality reduction offers RangeLDM a considerable advantage in terms of generation speed. As indicated in Table 4, the proposed method significantly outperforms LiDARGen [94] and UltraLiDAR [78] in terms of generation speed. In particular, our model is 200 times faster than LiDARGen and 30 times faster than UltraLiDAR.
@@
Method Years $\text{MAE}_{\text{range}}$downarrow $\text{Accuracy}^{\dagger}$uparrow $\text{IoU}^{\dagger}$uparrow
@@
Table 6: LiDAR upsampling. †: RangeNet++ semantic segmentation performance. PUNet, Grad-PU and DeepRS have no semantic segmentation performance as they do not upsample intensity.
5.2.2 Evaluation on the NuScenes Dataset
The results listed in Table 6 demonstrate that our model outperforms LiDAR VAE and LiDARGen significantly over nuScenes. In particular, our results on the MMD metric are an order of magnitude lower than those of other methods, while we reduce to about $1/3$ of the results of competitive methods in terms of the JSD metric. Also, Figure 6 illustrates the superiority of our model over LiDAR VAE and LiDARGen in terms of qualitative comparison.
5.3 Conditional Generation
LiDAR Point Cloud Upsampling.
Hough Voting Range-Guided Discriminator Circular Convolution Direction Conditioned $\text{MMD}_{\text{BEV}}$downarrow $\text{FRD}_{\text{range}}$downarrow $\text{JSD}_{\text{BEV}}$downarrow
(a) $\times$ $\times$ $\times$ $\times$ $3.95\times 10^{-4}$ $1536.7$ $0.067$
(b) $\checkmark$ $\times$ $\times$ $\times$ $6.57\times 10^{-5}$ 1229.3 0.056
Table 7: Main ablation study.
We obtain sparse input point clouds by selecting a subset of 16 beams from the raw 64-beam sensors, in line with LiDARGen. We compared our approach against PUNet [87], DeepRS [9], Grad-PU [23], bicubic interpolation, Nearest Neighbor (NN) interpolation and LiDARGen.
@@ 5.3 Conditional Generation
LiDAR Point Cloud Inpainting.
To mimic the case of missing data in LiDAR point clouds, we mask point clouds in front of the vehicle within a $22.5^{\circ}$ range and perform inpainting using LiDARGen and our method. Our results achieve much less MAE ($0.190$) between the reconstructed results and ground truth compared to that of LiDARGen ($0.367$). As shown in Figure 7, LiDARGen generates roads but fails to recover the masked car (the first row) and the wall (the second row), while we are able to reconstruct the details. More results are presented in the supplementary materials.
@@ 5.4 Ablation Study
Model VAE Model Size (M) $\text{MMD}_{\text{BEV}}$downarrow $\text{FRD}_{\text{range}}$downarrow $\text{JSD}_{\text{BEV}}$downarrow Throughput $\uparrow$ (samples/s)
@@
Discriminator $\text{PSNR}_{\text{range}}$uparrow $\text{MAE}_{\text{range}}$downarrow $\text{FRD}_{\text{range}}$downarrow CD $\downarrow$
@@
Table 9: Ablation studies on the VAE architecture.
As demonstrated in Table 7, we investigate the contribution of each component. Starting from the backbone model (a) without any component, we gradually add each component for evaluation. By applying Hough Voting to the training, the performance of variant (b) improves significantly, demonstrating the huge impact of the correct range-view data distribution. With the addition of the range-guided discriminator component, model (c) achieves better performance, thus clarifying the effectiveness of the geometry-sensitive module. Compared to model (c), the performance of model (d) decreases because circular convolution generates LiDAR scenes with arbitrary directions due to rotation invariance. However, when combined with direction-conditioned generation that guides the model to generate point clouds in the correct direction, the overall performance improves, as demonstrated by model (e). We refer the readers to the supplementary materials for qualitative results.
Latent diffusion.
We explore the contribution of latent diffusion to the generation quality and generation speed of the model. As shown in Table 8, we constructed two variants of different sizes that directly generate range images with diffusion models, denoted as “Ours-DMS” (with a smaller size) and “Ours-DM” (with a larger size), for comparison. It turns out that a much larger model size is required for “Ours-DM” to achieve competitive performance with “Ours”, due to the multitude of high-frequency details contained in the range image. Additionally, the superior throughput of “Ours” demonstrates the efficiency of latent diffusion.
VAE Architecture.
We investigate the contribution of the proposed range-guided discriminator by comparing our VAE with two baselines: 1) VAE without discriminator; and 2) VAE with vanilla discriminator (i.e., using the original 2D convolution in the discriminator). We evaluate the reconstruction performance of VAEs with 2D and 3D metrics, including $\text{PSNR}_{\text{range}}$, $\text{MAE}_{\text{range}}$, $\text{FRD}_{\text{range}}$ and Chamfer Distance (CD) [16]. As listed in Table 9, the VAE with the proposed range-guided discriminator outperforms the baselines in both 2D and 3D reconstruction metrics. This gives credits to the ability of perceiving local 3D structures by the range-guided discriminator. The qualitative results presented in Figure 8 also indicate that the proposed range-guided discriminator ensures the point clouds are reconstructed with less noise and more precise object structures.

### v2

@@ 4.4 Implementation Details
Circular Convolution. Range images are inherently circular, which means that the left boundary of a range image is actually connected to its right boundary. However, standard convolutions employ zero-padding or symmetry padding and do not take into account such constraints. Thus, as in LiDARGen [95], we replace standard convolutions in both the VAE and LDM with circular convolutions [59], which treat the left and right boundaries as connected neighbors in the topology. Since the 2D convolution operator satisfies translation invariance, it is evident that the entire network is invariant to horizontal shifting of the range image (i.e., rotation in the xy-plane of the point cloud).
@@
Conditioning the Model with Direction. Real LiDAR point clouds collected from driving scenes usually have a certain directionality. For example, since vehicles typically follow the direction of the road, the x-axis direction of the LiDAR point cloud usually aligns with the road direction. However, since the whole network of our model is invariant to rotation in the xy-plane of the point cloud, the point clouds generated from random noise could result in arbitrary directions. To avoid this, we generate point clouds with directional conditioning. Specifically, we use an $h\times w$ matrix as the condition, in which only the values from the first column are set to $1$, while the rest are all set to 00. This $\mathbf{1}$-valued column corresponds to the x-axis of the point cloud. Such simple conditioning demonstrates effective control over the direction of the generated point cloud, as presented in the qualitative results of the supplementary material.
Network Architectures and Model Hyperparameters. For simplicity, we adopted a similar architecture to that in [55]. The downsampling factor $f$ is set as $4$ to strike a good balance between efficiency and realistic results. The VAE comprises three encoder blocks and three decoder blocks. The LDM consists of four downsampling blocks with the last three followed by a transformer layer, and four upsampling blocks with the first three blocks also followed by a transformer layer. In the generation process, we employed 50 denoising steps with DDIM sampler [66] for point cloud generation. More implementation details can be found in released code https://github.com/WoodwindHu/RangeLDM.
@@ 5.1 Experimental Setup
Datasets. We evaluate our model on two challenging datasets, KITTI-360 and nuScenes. KITTI-360 has 81,106 LiDAR readings from 9 sequences in Germany, covering diverse scenes. We used the first two sequences for validation and trained on the rest. NuScenes is a public dataset with 297,737 LiDAR sweeps in the training set and 52,423 in the testing set, collected in Boston and Singapore.
Evaluation metrics. Following [95], , we employed three metrics to perform quantitative analysis: Maximum Mean Discrepancy (MMD), Jensen-Shannon divergence (JSD), and Frechet Range Distance (FRD score). We use a $100\times 100$ 2D histogram on the BEV plane to calculate MMD and JSD metrics. The Frechet Range Distance (FRD) score [95] is a metric used to evaluate the quality of samples acquired by a generative model, inspired by the FID score for images [24]. To compute the FRD score, we use RangeNet++ [45], an encoder-decoder-based network for segmentation, which is pre-trained on KITTI-360.
Baselines. We evaluated our approach for LiDAR point cloud generation against several competitive methods, including LiDAR VAE [5], LiDAR GAN [5], Projected GAN [58], LiDARGen [95] and UltraLiDAR [79].
5.2 Unconditional Generation
Method Years $\text{MMD}_{\text{BEV}}$downarrow FRD $\downarrow$ $\text{JSD}_{\text{BEV}}$downarrow
@@ 5.2 Unconditional Generation
Quantitative Results on KITTI-360. Table 2 displays the quantitative results of the proposed RangeLDM and competing algorithms. The results clearly illustrate that our method significantly outperforms the baselines across all metrics. Notably, our results on the MMD metric are an order of magnitude lower than those of other methods. This underscores the remarkable expressive capabilities of our model in the context of LiDAR point cloud generation.
Qualitative Results on KITTI-360. Figure 4 displays a set of randomly generated samples from competing algorithms, alongside real point cloud samples extracted from the KITTI-360 dataset for comparison. We observe that the Projected GAN [58] exhibits noticeable artifacts in the distant range. LiDARGen [95] effectively captures the general layout, but produces rather noisy outputs and lacks the same degree of straight walls as actual samples. UltraLiDAR [79] generates structured and reasonable scenes, yet it only creates voxelized point clouds (as magnified in the first row) without intensity features. In contrast, our model consistently surpasses the baseline models and yields results closely resembling real-world data. For example, we can generate cars (emphasized by red boxes in the first row of Figure 4) and road bollards (red boxes in the second row of Figure 4) on the road. Additionally, our model generates various scenes such as straightway (the first two rows of Figure 4) and crossroads (the last row of Figure 4). The supplementary materials provide more visualization results.
@@
Table 3: Human study on KITTI-360.
Human study on KITTI-360. To assess the perceptual quality, we conducted an A/B test involving a group of 14 researchers with LiDAR expertise. Following the same evaluation system as [95] and [79], we present pairs of randomly selected images from two point clouds and ask participants to determine which one appeared more realistic. The results, as displayed in Table 3, unequivocally demonstrate the superior visual quality of generation results by our model. In most of the cases, testers favored our results over the baselines.
@@
Evaluation on the NuScenes Dataset. The results listed in Table 4 demonstrate that our model outperforms LiDAR VAE and LiDARGen significantly over nuScenes. In particular, our results on the MMD metric are an order of magnitude lower than those of other methods, while we reduce to about $1/3$ of the results of competitive methods in terms of the JSD metric. Also, Figure 6 illustrates the superiority of our model over LiDAR VAE and LiDARGen in terms of qualitative comparison.
Model Size. The number of parameters in LiDARGen [95] and UltraLiDAR [79] are 29.7M and 40.3M, respectively. To ensure fair competition, we limited the capacity of our model to be similar to other approaches. Our model comprises 41.4M parameters, consisting of 12.7M parameters for VAE and 28.7M parameters for LDM. As illustrated in Table 2, our model outperforms the baselines by a large margin under similar model sizes.
Method Throughput $\uparrow$ (samples/s)
LiDARGen [95] 0.02
UltraLiDAR [79] 0.16
Ours 4.86
Table 5: Inference speed on a single RTX 3090 GPU.
Generation Efficiency. The proposed dimensionality reduction offers RangeLDM a considerable advantage in terms of generation speed. As indicated in Table 5, the proposed method significantly outperforms LiDARGen [95] and UltraLiDAR [79] in terms of generation speed. In particular, our model is 200 times faster than LiDARGen and 30 times faster than UltraLiDAR.
5.3 Conditional Generation
Hough Voting Range-Guided Discriminator Circular Convolution Direction Conditioned $\text{MMD}_{\text{BEV}}$downarrow FRD $\downarrow$ $\text{JSD}_{\text{BEV}}$downarrow
(a) $\times$ $\times$ $\times$ $\times$ $3.95\times 10^{-4}$ $1536.7$ $0.067$
(b) $\checkmark$ $\times$ $\times$ $\times$ $6.57\times 10^{-5}$ 1229.3 0.056
Table 6: Main ablation study.
LiDAR Point Cloud Upsampling. We obtain sparse input point clouds by selecting a subset of 16 beams from the raw 64-beam sensors, in line with LiDARGen. We compared our approach against PUNet [88], DeepRS [9], Grad-PU [23], bicubic interpolation, Nearest Neighbor (NN) interpolation and LiDARGen.
Method Years MAE $\downarrow$ Accuracy $\uparrow$ IoU $\uparrow$
@@ 5.3 Conditional Generation
Table 7: LiDAR upsampling.
@@
LiDAR Point Cloud Inpainting. To mimic the case of missing data in LiDAR point clouds, we mask point clouds in front of the vehicle within a $22.5^{\circ}$ range and perform inpainting using LiDARGen and our method. Our results achieve much less MAE ($0.190$) between the reconstructed results and ground truth compared to that of LiDARGen ($0.367$). As shown in Figure 7, LiDARGen generates roads but fails to recover the masked car (the first row) and the wall (the second row), while we are able to reconstruct the details. More results are presented in the supplementary materials.
@@ 5.4 Ablation Study
Model VAE Model Size (M) $\text{MMD}_{\text{BEV}}$downarrow FRD $\downarrow$ $\text{JSD}_{\text{BEV}}$downarrow Throughput $\uparrow$ (samples/s)
@@
Main Ablation. As demonstrated in Table 6, we investigate the contribution of each component. Starting from the backbone model (a) without any component, we gradually add each component for evaluation. By applying Hough Voting to the training, the performance of variant (b) improves significantly, demonstrating the huge impact of the correct range-view data distribution. With the addition of the range-guided discriminator component, model (c) achieves better performance, thus clarifying the effectiveness of the geometry-sensitive module. Compared to model (c), the performance of model (d) decreases because circular convolution generates LiDAR scenes with arbitrary directions due to rotation invariance. However, when combined with direction-conditioned generation that guides the model to generate point clouds in the correct direction, the overall performance improves, as demonstrated by model (e). We refer the readers to the supplementary materials for qualitative results.
Latent diffusion. We explore the contribution of latent diffusion to the generation quality and generation speed of the model. As shown in Table 8, we constructed two variants of different sizes that directly generate range images with diffusion models, denoted as “Ours-DMS” (with a smaller size) and “Ours-DM” (with a larger size), for comparison. It turns out that a much larger model size is required for “Ours-DM” to achieve competitive performance with “Ours”, due to the multitude of high-frequency details contained in the range image. Additionally, the superior throughput of “Ours” demonstrates the efficiency of latent diffusion.
Discriminator $\text{PSNR}_{\text{range}}$uparrow MAE $\downarrow$ FRD $\downarrow$ CD $\downarrow$
@@
Table 9: Ablation on the VAE architecture.
VAE Architecture. We investigate the contribution of the proposed range-guided discriminator by comparing our VAE with two baselines: 1) VAE without discriminator; and 2) VAE with vanilla discriminator (i.e., using the original 2D convolution in the discriminator). We evaluate the reconstruction performance of VAEs with 2D and 3D metrics, including PSNR, MAE, FRD and Chamfer Distance (CD) [16]. As listed in Table 9, the VAE with the proposed range-guided discriminator outperforms the baselines in both 2D and 3D reconstruction metrics. This gives credits to the ability of perceiving local 3D structures by the range-guided discriminator. The qualitative results presented in Figure 8 also indicate that the proposed range-guided discriminator ensures the point clouds are reconstructed with less noise and more precise object structures.